OpenStudy (lin.ivory):

Determine an expression equivalent to cos(Π+2A) a.-cos2A b.cos2A c.-sin2A d.sin2A

OpenStudy (ash2326):

@lin \(\cos x\) is negative in the third quadrant Therefore \(\large \cos (\pi+x)=-\cos x\)

OpenStudy (lin.ivory):

but dont i need to exapnd it with addition identities for these kind of questions?

OpenStudy (ash2326):

@lin.ivory

OpenStudy (ash2326):

Nope not needed here

OpenStudy (lin.ivory):

you saved me twice today!! Thanks. =)

OpenStudy (ash2326):

Welcome @lin.ivory

OpenStudy (anonymous):

You can do it with the addition identities if proof is required.

OpenStudy (ash2326):

\[\cos (\pi+2A)=\cos \pi cos 2A-\sin \pi \sin 2A\] \[\cos \pi=-1\] and \[\sin \pi=0\] \[\cos (\pi+2A)=-1 \times \cos2A-0=-\cos 2A\]

OpenStudy (lin.ivory):

but for trig equation questions, how do i determine whether what to use? theres the graphs, then the four quadrant ones, and then the identities, how do i know which one would give me the direct answer?

OpenStudy (ash2326):

Depending on points. If it's 1 point then just give the answer If it's more than that use the identities and get to the final result

OpenStudy (lin.ivory):

sorry for all the questions, but what is "1 point" though? you mean like cosx is 1 point and cos2x is two points?

OpenStudy (anonymous):

I think he means 1 mark is available for writing a correct answer.

OpenStudy (lin.ivory):

HAHA!! gotcha! thx @IceCreamVan so it just depend on whether i need to show my works or not?! THANKS! =)

OpenStudy (anonymous):

Yup! Always show your working when necessary though, one can never tell!

OpenStudy (lin.ivory):

=) of course!